Please answer question in bold. Thanks

Lesson 5 Roundup Exercise Assignment

In Lesson Five you've learned the basics of probability theory and how to calculate probabilities of outcomes in various circumstances.

  1. Provide a real-world example of classical probability and explain how the example illustrates the concept.
  2. Provide a real-world example of empirical probability and explain how the example illustrates the concept.
  3. Provide a real-world example of subjective probability and explain how the example illustrates the concept.
  4. Describe a situation where you could use the multiplication rule to determine the probability of an outcome. Include an example of the calculations.
  5. Describe a situation where you could use the addition rule to determine the probability of an outcome. Include an example of the calculations.
Answer #1

Answer:

By using given data,

classical probability:

It is the classical concept that mean measure the likelihood of somthing happening.in a classic sense it means that every statistical experiment will contain etimate that are equally likely happen.

This concept of classical probability .It the simplest form of probability that has equal adds to something happening.

Ex: This typical example of classical probability would be rolling a fair disc because it is equally probability that top of die will be an of the 6

no of this die :1,2,3,4,5,6

Emperical probability:

It is an event found through observation and experiment.it is likelihood that event will happen its based on this result of data collected.

Ex:Suppose we were surveya group of 50 chifs we ask them to choose their favourite type of food such as the italin,maxican,french,america or chinese choices.if 14 of the chef choose chinese as their fav food then based on these result the emperial probabilty that a cheif in this group would prefer chinese food over all other type of cuisine would be 0.28.this is an eg of actual abservation by collecting data through our suvery our probability is based on the result of this data.

Subjective probability:

subjective probability is where you use your opinion to find probability.

Ex:

you think you have an 80% chance of your best friend calling today because her broken down yesterday and she ill probabily need a ride.

Multiplication rule:

The multiplication rule is a way to find probability of two event happening at the same those.There are two multiplication rule P(AnB)=P(A).P(B) and specification rule is P(A) and P(B)=P(A)*P(B).P(A/B) means the probability of A happening that B has occured.

Additional rule:

When two events A and B are mutually exclusive the probability that A and B will occur is the sum of probability of each event.

P(A or B)=P(A)+P(B)